OpenStudy (anonymous):

Both y = 2e^t and y = e^t are solutions of the differential equation dy/dt = y. Since this equation is autonomous the graph of y=2e^t can be obtained from the graph of y=e^t by horizontal translation. By how much does one have to translate the graph of y=e^t horizontally to obtain the graph of y=2e^t? Does this question make sense?

OpenStudy (turingtest):

I'm thinking ln2\[\large e^{t+\ln2}=2e^t\]...makes sense to me at least :D

OpenStudy (anonymous):

so what would the horizontal translation be?

OpenStudy (turingtest):

ln2 units to the left

OpenStudy (turingtest):

we are changing\[f(t)\to f(t+\ln2)\]which is a translation of ln2 units left

OpenStudy (anonymous):

do you know what the answer to this is? Find the general solution to (b) dy/dt = t^2y^3 i think i have the right answer but want to be sure.

OpenStudy (turingtest):

what did you get?

OpenStudy (anonymous):

i got \[y=\pm \sqrt{3/2t^3 + c}\]

OpenStudy (turingtest):

I think you lost a minus sign, no?

OpenStudy (anonymous):

theres a very good chance .. what should the asnwer look like?

OpenStudy (turingtest):

\[\int\frac{dy}{y^3}=\int t^2dt\]\[-\frac1{2y^2}=\frac{t^3}3+C\to-\frac3{2t^3+C}=y^2\]\[y=\pm\sqrt{-\frac3{2t^3+C}}\]which would change the domain

OpenStudy (anonymous):

thank you!